OpenStudy (anonymous):

1. Solve the equation 4x^2 - 28x = 0. Check your solution(s) and state the final solution set.

OpenStudy (nikato):

What do 4x^2 and 28x have in common?

OpenStudy (anonymous):

The x? @nikato

OpenStudy (nikato):

Yes and???

OpenStudy (anonymous):

They each have a 7 in common, if you mean that @nikato I'm really bad with this stuff, if you couldn't tell haha

OpenStudy (solomonzelman):

\[4x^2 - 28x = 0\]\[4x(x - 7) = 0\]

OpenStudy (solomonzelman):

I am factoring it.

OpenStudy (nikato):

No.4 is not divisible by 7. If you list out all the factors 4: 1,2,4 28: 1,2,4,7,14,28 You should see that the largest factor they have in common is 4 So 4x^2 and 28x both have 4x in common

OpenStudy (anonymous):

Okay, I ended up with this 4x=0 4x/4=0/4 x=0 x-7=0 x-7+7=0+7 and my solutions were x=7 x=0

OpenStudy (haseeb96):

4x^2 -28x =0 okay now take 4 and x common from the above equation 4x(x-7)=0 now 4x(x-7)=0 either 4x=0 or x -7=0 x=0/4 or x=7 x=0 or x=7 this is your answer and solution of how to solve the equation

OpenStudy (anonymous):

Thank you @Haseeb96 and @nikato

OpenStudy (nikato):

No problem

OpenStudy (haseeb96):

welcome miss singer